Cove’s Cakes is a local bakery. Price and cost information follows:

Price per cake $ 14.21
Variable cost per cake
Ingredients 2.31
Direct labor 1.18
Overhead (box, etc.) 0.18
Fixed cost per month $ 4,637.60


Required:
1.
Determine Cove’s break-even point in units and sales dollars. (Round your Break-Even Units answer to the nearest whole number. Round your other intermediate calculations and sales dollars answer to 2 decimal places.)

   

2. Determine the bakery’s margin of safety if it currently sells 500 cakes per month. (Round your intermediate calculations to 2 decimals. Round the break-even units and final answer to nearest whole dollar.)

   

3. Determine the number of cakes that Cove must sell to generate $2,000 in profit. (Round your intermediate calculations to 2 decimal places and final answer to nearest whole number.)
Answer #1

Break even units = Fixed cost /Contribution margin per unit

= 4637.60/(14.21-2.31-1.18-0.18)

= 440 units

Breakeven even sales = 440 * 14.21

= 6,252.40

Margin of Safety = Sales units - Breakeven units

= 500 - 440

= 60

Sales needed = Fixed cost + desired profit /contribution margin per unit

= (4637.60+2000)/ (14.21-2.31-1.18-0.18)

= 6637.6/10.54

= 630